Q10 An article contained the following observations on degree of polymerization for paper specimens for which viscosity times concentration fell in a certain middle range: 420 425 427 427 432 433 434 437 439 446 447 448 453 454 465 469 Suppose the sample is from a normal population.

(a) Calculate a 95% confidence interval for the population mean, and interpret it.

(b) Calculate a 95% upper confidence bound for the population m
Answer #1

from the data we can easily calculate

sample mean =

sample SD=S=14.34

sample size=n=16

since we don't have population standard deviation so we will use t statistics which have df=n-1=16-1=15

a)

now 95% confidence interval is given by

so interval is (433.36,448.64)

b)

now 95% upper confidence bound is given by

so confidence bound is 447.28